OpenStudy (anonymous):

Solve log4 x – 1 = 2 Solve log4 x – 1 = 2 @Mathematics

OpenStudy (anonymous):

Is it 64

OpenStudy (anonymous):

yes, log4 x = 3 4^3 = x therefore x=64

OpenStudy (anonymous):

Thank you can you help me with another problem?

OpenStudy (anonymous):

quite possibly

OpenStudy (anonymous):

Simplify log5 s + log5 11 – log5 15 – log5 t

OpenStudy (lilg132):

no its not 64 its 17

OpenStudy (anonymous):

I put 17 before, but its not the correct answer

OpenStudy (anonymous):

Since the bases are the same apply the rules of logs to them. =log5 (11s) - log5 (15/t) =log5 (11*15s/t)

OpenStudy (anonymous):

But I get this s * 11 / 15 / t

OpenStudy (anonymous):

oh sorry so that will re arrange to be 11st/15 then

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

?